# Deep Feature-specific Imaging

### v5

@@ 4.2 Simulated Experiments Design
In this section, we present two distinct classification tasks. Initially, we evaluated the effectiveness of various coding schemes using the MNIST dataset, focusing on their performance in number recognition. Subsequently, we extend the evaluation to CIFAR10 by OViT , examining the adaptability and efficacy of DeepFSI in this different domain.
The MNIST dataset, which consists of handwritten digits ranging from 0 to 9, has a default size of 28 by 28 pixels [36]. To align the data with Hadamard matrices, we add black pixels to the edges of the images and resize them to 32 by 32 pixels [36]. Subsequently, we rescale the pixel values from the original range of $[0,1]$ to $[0.3,1]$. This rescaling introduces non-zero Poisson noise to the black regions, which has variance proportional to the expected photon counts.
@@ 4.3 Hardware Experiments Design
A key concern with DeepFSI is that its optimization relies on simulation, and its feasibility in real-world vision systems remains to be experimentally validated. We are also interested in whether DeepFSI is more robust than FSI when the experimental conditions do not perfectly match those in simulation. Therefore, we deploy the methodology implemented for the MNIST in Section 4.2 on a lab SPC. As shown in Fig. 4, the experiment utilizes a DMD (digital micromirror device), specifically the DLP 7000 model from Texas Instruments, and a PMT (photomultiplier tube) from PicoQuant, specifically the PMA Series. The training stage is conducted on a computer, and the parameters are deployed on the SPC to recognize handwritten number pictures displayed on a monitor. The testset consisted of ten pictures and each handwritten number appeared exactly once.
Similar to OViT experiments introduced in Section. 4.2, the optimal compression ratio is not available. We decide to use 0.09 in this case as its overall performance in the simulation is satisfactory. In other word TH, FSI and DeepFSI use 92 masks, while others use 1024 masks. All tests are repeated for three times, and the final performance is the average of the three trials. For each coding design, we test its performance when the total exposure times are 0.001, 0.01, 0.1, and 1.0 seconds.
@@ 5.1 Simulated Experiments on MNIST
In Fig. 7(b), we see that PN degrades the performance of any coding designs, and using non-FSI coding cannot provide a performance gain compared to RS. FSI presents better performance than non-FSI methods, although it is not designed for PN, which is consistent with the discovery of Nahalanobis et al. [10]. It highlights the necessity of using data priors under PN in a vision system. More significantly, DeepFSI outperforms traditional FSI, showing the advantage of using end-to-end optimization compared to PCA under PN. Additionally, we present a control group that first optimizes the coding under AGN and subsequently optimizes the classifier with this coding fixed under PN. We observe a suboptimal performance on the control group, indicating that using the optimal coding under AGN cannot lead to optimal performance in a vision system. It is an important discovery that the crucial role of noise models for the regime of future camera design.
5.2 Robustness Analysis
@@
Focusing at the values when number of photon is $10^{5}$, we can see a tremendous performance degradation in FSI if we choose an improper number of masks, although its performance supreme is much better than that of non-FSI methods, i.e. the truncated hadamard in this example. However, while the overall performance is no worse than FSI, DeepFSI presents more robust classification rates when the number of masks is suboptimal. Given that it is challenging to know the proper number of masks in real-world applications as it depends on the task and training data, DeepFSI has a significant feature of greater tolerance, which greatly improves the practicality compared to FSI.
@@
Figure 8: Classification rates on simulated data regarding compression ratios and light levels with (8(a)) Truncated Hadamard, (8(b)) FSI, and (8(c)) DeepFSI. The models are trained under Poisson noise using a light level of $10^{5}$ photons and subsequently tested under different light levels. It is observed that FSI performance deteriorates when the compression ratio is not properly chosen, whereas the Truncated Hadamard and DeepFSI models exhibit greater robustness in such scenarios.
@@ 5.3 OViT Simulations
Fig. 9 shows the classification rates on the CIFAR10 dataset with PN, based on simulation results. In this experiment, we assume the best compression rate is unavailable, and all coding designs employs $m=N$ masks where $N$ is the number of pixels in a patch. When PN is present, OViT outperforms all non-FSI coding strategies. This result shows DeepFSI’s feasibility in in more complicated vision tasks, and verifies its robustness with wrong compression ratio in Section 5.2, suggesting DeepFSI is a more practical method than FSI in general. Notably, our optical scanner module is simply a fully connected layer, much smaller in size compared to the ViT classifier. This further underscores that any software cannot ever fully compensate for the performance limitations of suboptimal hardware. This supports our perspective that once data are measured with suboptimal coding in any vision system, its overall performances on subsequent computer vision tasks are strictly bounded. All modern computer vision algorithms, unless they also optimize the hardware, are constrained by the upper bound set by the hardware and cannot exceed it, preventing them from fully exploiting their true performance potential.
@@ 5.4 Hardware Experiments on MNIST
Fig. 10 compares the classification rates of different coding designs in a hardware experiment. We see that DeepFSI outperforms the other coding by a huge margin. This phenomenon can be explained by the results in Section. 5.2 that DeepFSI provides better tolerance and performance than the other coding designs when the compression ratio used is not ideal or not available, which is fixed at 0.09 in this experiment. In real-world applications, noise can be more complex than what is considered in this work. Factors such as ambient light and dark counts can cause the vision system to deviate from its supreme achieved in training conditions, leading to unexpected performance degradation to other coding designs shown in Fig. 10. Therefore, the robustness and tolerance of DeepFSI are of great value in the promotion of task-specific imaging.
@@ 6 Discussion and Limitations
We propose a deep feature-specific imaging model to complement conventional feature-specific imaging techniques, which is originally designed for additive Gaussian noise, under Poisson noise. Its performances have been tested with a single-pixel camera compared to other coding designs. DeepFSI presents a better performance than that of FSI under Poisson noise and a better practicality and tolerance in real-world implementation as well. Since Poisson noise is a growing challenge for photon-counting technology in future vision systems design, the proposed model extended the applicability of sensors using such technology.
@@
In addition to some facts already known to many, such as that a higher light throughput may not guarantee a better performance which is shown in Fig. 7, this work also discloses several other interesting phenomena. First, using AGN-optimal coding under PN is not recommended as the computed gradient on optical coding layer in these two models are different, indicating different directions during optimization. Second, future camera or vision system design must incorporate hardware optimization to achieve the best performance. However, this end-to-end optimization must be performed under the correct noise model.
One limitation in this work is that we use the MLGauss noise model whose variance and mean are equal and apply reparameterization to approximate the actual photon noise at the sensor during training. This approach offers a valid alternative to Poisson noise in gradient estimation when the number of photons is greater than 10 [13], but it may not provide the same performance gain at extremely low-light conditions, where the actual Poisson noise for testing deviates more from it. A potential future direction is to develop a more accurate reparameterization model tailored for such low-light scenarios.

### v6

@@ 4.2 Simulated Experiments Design
In this section, we present three classification tasks to explore the upper bound performance of the proposed framework under various hyperparameter configurations, following a standard vision system implementation workflow that includes simulation-based optimization and real-world deployment. The hardware tests, presented in the following section, evaluate the transfer robustness of the optimized design as an integral part of this workflow. Initially, we evaluated the effectiveness of various coding schemes using the MNIST dataset, focusing on their performance in number recognition. Subsequently, we extend the evaluation to CIFAR10 by OViT , examining the adaptability and efficacy of DeepFSI in this different domain. We also test the proposed method on hyperspectral data to see its performance on more realistic scenarios. The MNIST dataset, which consists of handwritten digits ranging from 0 to 9, has a default size of 28 by 28 pixels [37]. To align the data with Hadamard matrices, we add black pixels to the edges of the images and resize them to 32 by 32 pixels [37]. Subsequently, we rescale the pixel values from the original range of $[0,1]$ to $[0.3,1]$. This rescaling introduces non-zero Poisson noise to the black regions, which has variance proportional to the expected photon counts.
@@
To further evaluate the generalization capability of DeepFSI beyond conventional image classification benchmarks, we conduct an additional experiment on hyperspectral data, which better reflects real-world sensing applications. In this experiment, we utilize the Indian Pines dataset [39], which was acquired through the Airborne Visible/Infrared Imaging Spectrometer (AVIRIS) on June 12, 1992, covering the Purdue University Agronomy farm northwest of West Lafayette and its surrounding area. This dataset consists of $145\times 145$ pixels with 224 spectral reflectance bands spanning wavelengths from 0.4 to 2.5 $\mu m$, providing rich spectral information for material classification. The goal is to assign each pixel’s spectral signature to a material category. Fig. 4(a) illustrates a representative spectral band (band 12), along with the class distribution within the dataset and the corresponding ground-truth semantic labels. Fig. 4(b) further shows the pixel-wise spectral responses for two representative classes across the 224 channels. Unlike standard image classification, feature extraction in this experiment is performed in the spectral (wavelength) domain rather than the spatial domain. Spectral channels are zero-padded from 224 to 256 to align with the Hadamard mask and rescaled to the range $[0,1]$. We utilize the scanner-classifier module from Fig. 2, adapting the DeepFSI classification layers to the new dataset sizes. Each classification run for 2000 epochs, with a learning rate of $5\times 10^{-3}$ and a batch size of 5000. Classification performances are evaluated cross varying photon budgets from $10^{1}$ to $10^{10}$.
(a) Indian Pines dataset [39]
(b) Visualization of hyperspectrum examples
Figure 4: IndianPines dataset overview
@@ 4.3 Hardware Experiments Design
A key concern with DeepFSI is that its optimization relies on simulation, and its feasibility in real-world vision systems remains to be experimentally validated. In the simulated experiments of Section 4.2, the objective is to explore the theoretical performance upper bound of the proposed framework across a wide range of hyperparameters under controlled and idealized conditions. However, practical deployment inevitably introduces deviations from simulation assumptions. Therefore, the primary goal of the hardware experiment is not to further improve peak accuracy, but to evaluate the transfer robustness of DeepFSI from simulation to a physical system. In particular, we investigate whether DeepFSI remains robust and generalized compared to other methods when experimental conditions, such as photon budget, optimal compression ratio, device bit depth, and ambient illumination, differ from those assumed during simulation and may be partially unpredictable. To this end, we deploy the methodology developed for MNIST in Section 4.2 on a laboratory SPC system. As shown in Fig. 5, the setup consists of a DMD (digital micromirror device), specifically the DLP7000 model from Texas Instruments, and a PMT (photomultiplier tube), specifically the PMA Series from PicoQuant. The training stage is conducted on a computer, and the parameters are deployed on the SPC to recognize handwritten digit pictures displayed on a monitor. The testset consisted of ten pictures and each handwritten number appeared exactly once , but each digit is independently sampled at least 10 times to compute the average classification rates.
Similar to OViT experiments introduced in Section 4.2, the optimal compression ratio is not available. We decide to use 0.09 in this case as its overall performance in the simulation is satisfactory. In other word TH, FSI and DeepFSI use 92 masks, while others use 1024 masks. For each coding design, we test its performance when the total exposure times are 0.0001, 0.001, 0.01, and 0.1 seconds.
@@ 5.1 Simulated Experiments on MNIST
In Fig. 8(b), we see that PN degrades the performance of any coding designs, and using non-FSI coding cannot provide a performance gain compared to RS. FSI presents better performance than non-FSI methods, although it is not designed for PN, which is consistent with the discovery of Mahalanobis et al. [10]. It highlights the necessity of using data priors under PN in a vision system. More significantly, DeepFSI outperforms traditional FSI, showing the advantage of using end-to-end optimization compared to PCA under PN. Additionally, we present a control group that first optimizes the coding under AGN and subsequently optimizes the classifier with this coding fixed under PN. We observe a suboptimal performance on the control group, indicating that using the optimal coding under AGN cannot lead to optimal performance in a vision system. It is an important discovery that the crucial role of noise models for the regime of future camera design.
Photons Gaussian Poisson Value Distributions
$10^{3}$ Refer to caption Refer to caption Refer to caption
$10^{5}$ Refer to caption Refer to caption Refer to caption
$10^{8}$ Refer to caption Refer to caption Refer to caption
Figure 9: Masks optimized by DeepFSI on the MNIST dataset are shown. Each row displays six of the optimized masks under Gaussian and Poisson noise, along with their value distributions for a given number of photons. The intensity values range continuously from black to white, with black pixels near 1-1, gray pixels at 0, and white pixels near 11.
Fig. 9 illustrates several DeepFSI-optimized masks for the MNIST dataset. At low photon budgets, the mask value distributions for AGN and Poisson noise differ significantly, with Poisson masks focusing on the essential regions and performing repeated measurements to improves the SNR of those informative features, while maintaining lower light throughput compared to Gaussian masks. However, as the photon count increases, both masks converge to a similar configuration. This observation is consistent with the conclusions of Raginsky et al. [6], demonstrating that DeepFSI inherently learns to mitigate the SNR dilution effect characteristic of photon-starved sensing
5.2 Error and Robustness Analysis
5.2.1 Compression Robustness
@@
Focusing at the values when number of photon is $10^{5}$, we can see a tremendous performance degradation in FSI if we choose an improper number of masks, although its maximum performance is much better than that of non-FSI methods, i.e. the Truncated Hadamard in this example. However, while the overall performance is no worse than FSI, DeepFSI presents more robust classification rates when the number of masks is suboptimal. Given that it is challenging to know the appropriate number of masks in real-world applications as it depends on the task and training data, DeepFSI has a significant feature of greater tolerance, which greatly improves the practicality compared to FSI.
@@
Figure 10: Classification rates on simulated data regarding compression ratios and photon budgets with (10(a)) Truncated Hadamard, (10(b)) FSI, and (10(c)) DeepFSI. The models are trained under Poisson noise using a photon budget of $10^{5}$ and subsequently tested under different photon budgets. It is observed that FSI performance deteriorates when the compression ratio is not properly chosen, whereas the Truncated Hadamard and DeepFSI models exhibit greater robustness in such scenarios.
5.2.2 MLGauss Approximation
Since sampling under PN is not directly differentiable, we adopt the MLGauss method during training to enable gradient-based optimization. To assess the effect of this approximation, we evaluate DeepFSI under both MLGauss and PN models.
Fig. 11 compares performance under matched conditions (MLGauss-training-with-MLGauss-testing) and mismatched conditions (MLGauss-training-with-Poisson-testing). Noticeable discrepancies between the two testing models mainly arise in the extremely low-photon regime (below 0.1 photons per pixel per measurement), where the overall photon budget is highly limited and classification accuracy is substantially reduced. Outside this regime, the difference between MLGauss and Poisson testing remains small, suggesting that MLGauss serves as a reasonable and practical approximation to Poisson noise in our setting.
(a) MLGauss training + MLGauss testing
(b) MLGauss training + PN testing
Figure 11: MLGauss analysis is presented in the contour plots above. Although some degradation is observed when MLGauss is used to train the model for applications under PN, the differences are primarily concentrated in the low light region, where the illumination is insufficient for meaningful classification.
5.2.3 Quantization
We also analyzed the effect of quantization error to evaluate the transfer of the simulated model to practical implementation on a DMD. In our system, the masks are quantized to the instrument accuracy of $1/768$, making it important to assess whether this level of quantization affects performance.
Fig. 12 compares DeepFSI performance with and without quantization using the MLGauss-training-with-Poisson-testing mode. The results show that quantization has minimal impact on classification accuracy, suggesting that the model is robust to the DMD’s discretization limits.
(a) No quantization under PN
(b) Quantization under PN
Figure 12: Quantization error analysis is performed by comparing the classification rates on DeepFSI with and without quantization, as shown in the contour plots above. The difference between the two modes is not readily apparent.
@@ 5.3 OViT Simulations
Fig. 13 shows the classification rates on the CIFAR10 dataset with PN, based on simulation results. In this experiment, we assume the best compression rate is unavailable, and all coding designs employs $m=N$ masks where $N$ is the number of pixels in a patch. When PN is present, OViT outperforms all non-supervised coding strategies. This result shows the feasibility of DeepFSI in more complex vision tasks. It also verifies its robustness to compression ratio mismatch, as analyzed in Section 5.2, suggesting that DeepFSI is generally more practical than FSI. Notably, our optical scanner module is simply a fully connected layer, much smaller in size compared to the ViT classifier. This further underscores that any software cannot ever fully compensate for the performance limitations of suboptimal hardware. This supports our perspective that once data are measured with suboptimal coding in any vision system, its overall performances on subsequent computer vision tasks are strictly bounded. All modern computer vision algorithms, unless they also optimize the hardware, are constrained by the upper bound set by the hardware and cannot exceed it, preventing them from fully exploiting their true performance potential.
@@
(a) $10^{2}$ photons
(b) $10^{4}$ photons
(c) $10^{6}$ photons
Figure 14: Visualization of masks optimized by OViT. Image patches are processed by the 8 by 8 masks sequentially.
Fig. 14 shows the OViT-optimized masks with different photon budgets. The DeepFSI model always prioritizes the most informative region in signals, and remains frugal in photon counting by reducing the light throughput. When the photon budget is constrained, DeepFSI refines the quality of selected features through repeated measurements with similar masks, rather than acquiring additional features. This behavior is consistent with the pattern previously shown in Fig. 9.
@@ 5.4 Hardware Experiments on MNIST
Fig. 15 compares the classification rates of different coding designs in a hardware experiment. The digit-specific classification performances are illustrated in Fig. 16. DeepFSI clearly outperforms the other coding designs in all digits when photon budget is sufficient. This phenomenon can be explained by the results in Section 5.2 that DeepFSI provides better tolerance and performance than the other coding designs when the compression ratio used is not ideal or not available, which is fixed at 0.09 in this experiment. In real-world applications, noise can be more complex than what is considered in this work. Factors such as ambient light and dark counts can cause the vision system to deviate from its best performances achieved in training conditions, leading to unexpected performance degradation to other coding designs shown in Fig. 15. Unfortunately, these factors cannot be fully addressed in the system design, and even achieving the precise number of photons required for coding optimization is rarely possible. Therefore, the robustness and tolerance of DeepFSI are of great value in the promotion of task-specific imaging.
@@
(a) Exposure time [0.1 sec]
(b) Exposure time [0.01 sec]
(c) Exposure time [0.001 sec]
(d) Exposure time [0.0001 sec]
Figure 16: Digit-wise classification rates. Each square shows the classification probability on corresponding digit.
To strictly evaluate the robustness of our method and avoid pseudoreplication, we performed statistical analysis at the image (cluster) level rather than treating individual repetitions as independent samples. Confidence intervals (95% CI) were calculated using cluster bootstrap resampling ($N=10$ digit classes), and statistical significance was determined using double-sided Wilcoxon signed-rank test to account for class-specific difficulty variations in table 1,2, and 3. The results suggest that DeepFSI achieves competitive and generally stable performance across digit classes and noise levels. Compared to other methods, DeepFSI exhibits smaller performance variations, indicating improved tolerance to mismatches between simulation and experimental conditions. While these findings do not constitute definitive statistical proof due to the sample size, they provide supportive evidence of the DeepFSI’s transferability and robustness in practical settings.
Table 1: Confience Interval by Cluster Bootstrap (mean, [lower, higher])
Exposure [sec] DeepFSI FSI Truncated Hadamard Raster Hadamard
0.1 $0.970,[0.940,1.000]$ $0.610,[0.420,0.800]$ $0.550,[0.350,0.740]$ $0.280,[0.000,0.560]$ $0.250,[0.090,0.440]$
0.01 $0.860,[0.720,0.960]$ $0.430,[0.300,0.560]$ $0.370,[0.200,0.560]$ $0.210,[0.030,0.420]$ $0.250,[0.080,0.440]$
0.001 $0.670,[0.520,0.810]$ $0.270,[0.150,0.410]$ $0.400,[0.260,0.540]$ $0.140,[0.050,0.260]$ $0.230,[0.110,0.360]$
0.0001 $0.327,[0.258,0.405]$ $0.160,[0.117,0.213]$ $0.310,[0.190,0.430]$ $0.107,[0.057,0.168]$ $0.130,[0.010,0.300]$
Table 2: Confidence Interval by Studentized Bootstrap (mean, [lower, higher])
Exposure [sec] DeepFSI FSI Truncated Hadamard Raster Hadamard
0.1 $0.970,[0.939,0.998]$ $0.610,[0.337,0.830]$ $0.550,[0.295,0.785]$ $0.280,[0.018,0.582]$ $0.250,[0.088,0.737]$
0.01 $0.860,[0.515,0.960]$ $0.430,[0.278,0.597]$ $0.370,[0.192,0.705]$ $0.210,[0.027,0.725]$ $0.250,[0.055,0.559]$
0.001 $0.670,[0.477,0.823]$ $0.270,[0.151,0.509]$ $0.400,[0.252,0.595]$ $0.140,[0.041,0.358]$ $0.230,[0.091,0.402]$
0.0001 $0.327,[0.253,0.434]$ $0.160,[0.115,0.291]$ $0.310,[0.156,0.446]$ $0.107,[0.054,0.233]$ $0.130,[0.012,0.793]$
Table 3: Significance Analysis VS DeepFSI (Double-sided Wilcoxon statistic, [p value], rank‑biserial)
Exposure [sec] FSI Truncated Hadamard Raster Hadamard
0.1 $1.000,[\mathbf{0.017}],0.944$ $0.000,[\mathbf{0.012}],1.000$ $1.500,[\mathbf{0.010}],0.933$ $0.000,[\mathbf{0.002}],1.000$
0.01 $0,[\mathbf{0.002}],1.000 0.000,[0.0$$0,[\mathbf{0.012}],1.000$ $1.000,[\mathbf{0.004}],0.964$ $0.000,[\mathbf{0.002}],1.000$
0.001 $3.000,[\mathbf{0.010}],0.891$ $5.500,[\mathbf{0.027}],0.800$ $0.000,[\mathbf{0.008}],1.000$ $2.500,[\mathbf{0.010}],0.909$
0.0001 $2.000,[\mathbf{0.006}],0.927$ $24.000,[0.770],0.127$ $0.000,[\mathbf{0.002}],1.000$ $9.000,[0.064],0.673$
5.5 Experiments on Indian Pines
(a) Classification under AGN
(b) Classification under PN
[…]
Fig. 17 illustrates the classification performances on Indian Pines. The overall trends are consistent with our observations on MNIST, demonstrating the generalization capability of DeepFSI to hyperspectral data. Although DeepFSI does not exhibit clear advantages under AGN assumption, its performance under PN assumption surpasses that of all other methods designed based on the AGN assumption. These results indicate that DeepFSI could maintain robust performance in higher-dimensional, more realistic sensing scenarios, highlighting its potential for generalization beyond conventional image classification benchmarks.
@@ 6 Discussion and Limitations
We propose a deep feature-specific imaging model to complement conventional feature-specific imaging techniques, which is originally designed for additive Gaussian noise, under Poisson noise. Its performances have been tested with a single-pixel camera compared to other coding designs. Instead of prioritizing pixel-level reconstruction, DeepFSI focuses on preserving task-relevant information and functional feature fidelity, enabling reliable task performance even when photon levels are insufficient for accurate image reconstruction. DeepFSI presents a better performance in classification tasks than that of FSI under Poisson noise and a better practicality and tolerance in real-world implementation as well. Since Poisson noise is a growing challenge for photon-counting technology in future vision systems design, the proposed model extended the applicability of sensors using such technology.
@@
In addition to some facts already known to many, such as that a higher light throughput may not guarantee a better performance which is shown in Fig. 8, this work also discloses several other interesting phenomena. First, using AGN-optimal coding under PN is not recommended as the computed gradient on optical coding layer in these two models are different, indicating different directions during optimization. Second, future camera or vision system design must incorporate hardware optimization to achieve the best performance. However, this end-to-end optimization must be performed under the correct noise model. Our results further suggest that in photon-limited regimes, the achievable performance of an FSI system is largely constrained by the efficiency of optical sampling at the front end, rather than solely by the complexity of the backend model. Optimizing the physical coding layer therefore enables more effective utilization of the available photons, beyond what can be achieved through digital post-processing alone.
The training complexity of DeepFSI is task specific. For DeepFSI on MNIST, the complete hyperparameter sweep (10 compression ratios, 4 photon budgets, 2 noise models, and 5 trials) required approximately $2.87\times 10^{4}$ seconds on an RTX 4500 Ada Generation GPU. Here, a single run refers to training and evaluating one specific combination of compression ratio, photon budget, noise model, and trial, which required approximately 71 seconds. For the hardware experiments, the primary time bottleneck lies in photon timestamp parsing rather than model inference. DeepFSI and FSI processing is relatively efficient, requiring approximately 1 second per run on an AMD Ryzen Threadripper PRO 7965WX (24-core) CPU, including parsing and final recognition. In contrast, Hadamard and Truncated Hadamard patterns are significantly more time-consuming due to their higher light throughput, requiring approximately 4 hours for 100 runs. Training OViT is more computationally demanding, requiring approximately 2.5 hours per run on a Quadro P5000 GPU. The batch size is selected to maximize GPU memory utilization. The influence of training data size on performance has not been systematically evaluated in this work and is left for future investigation.
One limitation in this work is that we use the MLGauss noise model whose variance and mean are equal and apply reparameterization to approximate the actual photon noise at the sensor during training. This approach offers a valid alternative to Poisson noise in gradient estimation when the number of photons is greater than 10 [13], but it may not provide the same performance gain at extremely low-light conditions, where the actual Poisson noise for testing deviates more from it. In our experimental settings, most methods that use a coding matrix generally operate well once the photon counts are sufficient for classification tasks (with more than one photon from each pixel before coding for each readout), as they combine photons from multiple pixels and exceed this threshold. Raster scan, however, may be more affected because it measures only a single pixel at a time. If the photon count falls below the level required for meaningful classification, errors in gradient estimation become noticeable, but in this regime they are a secondary concern since accurate classification is not achievable anyway. While we have observed this behavior for DeepFSI as shown in Fig. 11, other coding strategies may show similar trends, though further evaluation would be needed to confirm this.